# Quantifying artificial intelligence through algebraic generalization

### v1

@@ Introduction
Systematic compositional generalization and regression. Systematic compositional generalization refers to the ability to recombine known basis elements into novel combinations of fixed sequence size (Fig 4A). This means that test sets of systematicity are limited to novel combinations of the same length as seen during training Lake and Baroni, 2018, Hupkes et al., 2020. The analog in algebraic circuits is to 1) sample a family of circuits $\mathcal{C}$ over a field $\mathbb{F}$ of a fixed size and depth, and 2) to generalize to circuits of the same size and depth, but differentially sampling input gates and/or operators. This can be implemented by choosing different samplers – $P_{1}$ and $P_{2}$ – that differentially sample gates (Fig 4A). For example, the training set of circuits could be constructed with input gates $a_{i}\in_{P_{1}}\mathbb{F}$, and the testing set of circuits could be constructed with the input gates $b_{i}\in_{P_{2}}\mathbb{F}$. (Note that when samplers choose $b_{i}>>a_{i}$, this is analogous to the common test of “length generalization”; Fig. 2A.)
Interestingly, we note that successful systematic compositional generalization on input gates in algebraic circuit evaluations amounts to learning a distributionally robust regression model (e.g., Zhang et al. 2022, Ghosh et al. 2021). This is because a circuit of specific size and depth (e.g., as shown in Fig. 4A) expresses to a specific polynomial equation. Generalization on a circuit class to a distribution shift of its input gates (e.g., field elements but not operator gates) would demonstrate successful out-of-distribution generalization.
@@
Figure 4: Analogs of common compositional generalization benchmarks in terms of algebraic circuits. A) A simpler form of generalization within an algebraic circuit is generalizing to circuits of the same structure (size and depth), but with a novel combination of gates. The analog of this in compositionality is commonly-referred to as systematic compositional generalization Hupkes et al., 2020. (Learning over circuits of fixed structure can also be viewed as learning a regression model Zhang et al., 2022.) On the left, we illustrate an example of a model that is trained on a restricted family of circuits where the input gates are sampled from a field $\mathbb{F}$ with a sampling function $P_{1}$. At test time, the model is required to generalize to circuits of the same circuit class (in terms of size and depth), but where the input gates are sampled using a different sampler $P_{2}$. On the right, both input and operator gates are chosen with separate samplers across train and test circuits, resulting in a more difficult test of systematic compositionality. B) Productive compositional generalization considers partitions of the training and testing sets across circuit sizes and depths. The analog of this in compositionality is commonly-referred to as productive compositional generalization. Given a family of circuits $\mathcal{C}_{\mathbb{F},X,D,S}$, where $D$ and $S$ denote a set of depths $d_{i}$ and size $s_{i}$, the primary experimental manipulation is to construct a training set $\mathcal{C}_{\mathbb{F},X,D_{1},S_{1}}$ and testing set $\mathcal{C}_{\mathbb{F},X,D_{2},S_{2}}$ such that there exists no overlap of a specific circuit class $C_{\mathbb{F},X,d_{i},s_{i}}$ between the training and testing sets.
@@
Productive compositional generalization. Productive compositional generalization refers to the ability to generalize to sequences of longer length. In the context of algebraic circuits, while systematicity focuses on keeping circuit size and depth fixed while manipulating gates, productive generalization focuses on manipulating circuit size and depth (Fig 4B). Thus, evaluating systematic and productive generalization can be studied together; measures of generalization can be quantified in a continuous manner by varying circuit parameters, such as gate samplers (systematicity), size (productivity), and depth (productivity). Understanding how each of these properties interact across training and testing sets will provide a comprehensive quantification of algebraic generalization.
@@
Since algebraic circuits are circuit representations of algebraic expressions, one can ask more generic questions about algebraic polynomials. For example: Given a class of polynomials as a training dataset, what other class of polynomials will this model be able to compute? Such a question goes beyond asking whether a model can systematically or productively generalize. Instead, it addresses a basic question that can leverage other rich mathematical subfields (e.g., geometry, topology) to quantify AI generalization. Algebraic circuits provides a flexible framework to formalize problem complexity beyond existing paradigms in compositional generalization.
@@
Prior work studying arithmetic abilities in transformer models have typically focused on computing simple expressions with field elements ($a\in\mathbb{F}$; e.g., $5+7=?$), rather than variables ($x\in X$; e.g., $2x_{1}+x_{2}+7$). Including variables in an algebraic expression increases the polynomial degree ($d\geq 1$), thereby increasing its complexity and the need for abstractions. One approach to evaluating algebraic circuits with variables is through polynomial identity testing, an active area of research in computational complexity theory and computational algebra. Polynomial identity testing evaluates whether two polynomials are equivalent (i.e., $P_{1}(x_{1},...,x_{n})\equiv P_{2}(x_{1},...,x_{n})$). Posed another way, one can ask whether $P_{1}(x_{1},...,x_{n})-P_{2}(x_{1},...,x_{n})\equiv 0$.
@@
This problem can be naturally formulated as a binary classification task for AI models. Importantly, studying the extent of algebraic generalization by assessing train/test circuit divergences of polynomial identity problems can be rigorously quantified, since each identity problem can be encoded as a circuit (Fig. 5A). The complexity of a polynomial identity testing problem can be scaled up by increasing the size, depth, and/or degree of the test problems. Moreover, machine learning practitioners can leverage existing symbolic programs and solvers to systematically generate large datasets (e.g., SymPy; Meurer et al. 2017). Polynomial identity testing thus provides a unique opportunity to study AI generalization in terms of circuit complexity theory.
@@
Figure 5: Algebraic problems as machine learning challenges. Prior work that leverages arithmetic problems for machine learning studies are typically limited to evaluating expressions with field elements. We introduce problems that can be evaluated with abstract variables. A) Polynomial identity testing as a machine learning classification task. Polynomial identity testing, an important and active area of research in computational algebra, evaluates whether two polynomials are equivalent. We illustrate two different polynomial expressions, $P_{1}(x)$ and $P_{2}(x)$ with distinct circuit representations, yet are mathematically equivalent. To reformulate this as a classification task for machine learning studies, one can ask if $P_{2}(x)-P_{1}(x)\equiv 0$ (right). B) Polynomial expansion and/or factorization as a transduction (sequence-to-sequence) task. Common sequence-to-sequence tasks in linguistics ask whether a model can expand a string using a set of rules or a grammar, such as SCAN or PCFG Lake and Baroni, 2018, Hupkes et al., 2020. In algebraic circuits, this is analogous to expanding a polynomial in a factorized representation (left). Given the 1-1 correspondence of polynomials, an additional approach is to take a polynomial in its expanded form (sum of monomials), and generate the factorized representation (right). This provides the ability to evaluate whether an AI system can expand an encoding (expansion), or compress an encoding (factorization).
@@
A wide application of generative models is in sequence-to-sequence transduction tasks. One particularly impactful use case is the development of AI models for code. AI models for code take in code as inputs (e.g., COBOL), and generate a translation of that code (e.g., Java). Despite its potential importance for modernizing many existing codebases, there is significant skepticism as to whether generative models trained with next-token prediction can reliably generate accurate code translations. This is due to the fact that programming languages are not dictated by autoregressive processes, and instead governed by symbolic rules.
Learning on algebraic circuits provides a straightforward framework to evaluate the ability of models to learn sequence-to-sequence tasks that are governed by symbolic rules. For example, the problem of expanding or factorizing a polynomial is a problem that is governed by the axiomatic rules of algebra (Fig. 5B). Like code translation, this task requires transforming one sequence into another while maintaining mathematical equivalence (or for code, algorithmic equivalence). Importantly, there are explicit tools to describe the complexity with which the translation occurs in algebraic circuits. For example, a polynomial expansion transforms a factorized representation (i.e., compact circuit) to a sum of monomials representation (i.e., larger circuit). In contrast, a polynomial factorization implements the inverse operation, which requires compressing a large circuit to a smaller circuit. While a general interesting question is to ask whether models trained on one type of transformation can learn the other, this formalization has natural implications for understanding how to design AI models for code. In particular, some programming languages may have lower-level syntax (e.g., COBOL) relative to other languages, such as Python or Java. In either case, understanding how lower-level and higher-level encodings of an expression (be they algebraic or programmatic) requires learning useful abstractions and the symbolic rules to translate them. Thus, studying the conditions by which AI models can robustly parse and translate algebraic circuits can shed light on the best strategies to train AI models across different programming languages.
@@
A major issue in assessing symbolic generalization is the difficulty of interpreting what goes awry when they fail to generalize. This is partly due to the lack of interpretability of many benchmarks, which are often presented in natural language, and where the ground truth algorithms (e.g., circuit diagrams or parse trees) are not known Clark et al., 2019, Hendrycks et al., 2020. A more recent set of approaches in characterizing the interpretability of neural network representations rely on the design of carefully constructed tasks that have interpretable task components, such as ground truth parse trees or circuit diagrams Andreas, 2018, Dziri et al., 2023, Deletang et al., 2022. Algebraic circuits similarly provide ground truth interpretability. While the input to an AI model might be a string representation of a polynomial, that polynomial’s circuit encoding provides its ground truth encoding. Such an encoding makes it easy to directly compare to the internal representations of a model, such as the attention weights within a transformer (Fig. 6). This makes it possible to ask if the internal representations of a transformer computes the algebraic expression similarly to its computable circuit encoding (the adjacency matrix of the graph) – the normative approach to computing polynomials. More broadly, the distance metric that computes the distance between attention weights and ground truth circuit edges (as shown in Fig. 6) can be used as a regularization term to encourage transformers to learn normative circuit computations, providing algorithmic and step-by-step information to the model. Overall, evaluating AI models on algebraic problems allows practitioners to adjudicate between theoretical models of circuit computation with modern AI computation.
Figure 6: Using an algebraic circuit’s adjacency matrix as a ground truth comparison to interpret transformer attention representations. The transformer’s attention mechanism provides a useful way to peer into the representations of its input tokens (Left). When the input is an algebraic expression (presented as a string of tokens), the attention matrix can be investigated to uncover the relationships between tokens (i.e., operators and operands). On the other hand, an algebraic expression can always be mapped to a ground truth circuit, which can be represented as the adjacency matrix of a circuit (Right). This allows for the direct comparison between attention weights and the ground truth circuit representation (Middle). This distance between the transformer’s attention weights and the ground truth circuit adjacency matrix can also be potentially used as a regularizer to encourage learning circuit representations in attention weights.
@@
We have introduced algebraic circuit tasks within the context of training models from scratch. While it is difficult to evaluate pretrained LLMs with precise levels of certainty due to the obscure nature of the pretraining data and optimization protocols, algebraic tasks can be still repurposed for LLMs. Prior work has demonstrated that LLMs typically fail on mathematical problems which require abstract reasoning on variables Nezhurina et al., 2024, McCoy et al., 2023, Wu et al., 2024. Thus, if an LLM calls specific tools (e.g., calculators) which make computing circuits with only field elements too simple (e.g., those illustrated in Fig. 1 and 2, we can replace these with circuits containing variables $x\in X$ (i.e., those of degree $\geq 1$)). For example, such tasks can include the aforementioned polynomial identity testing or polynomial expansion/factorization tasks, both of which can be made arbitrarily difficult and use arbitrary sets of variables/tokens.
Given that prior studies have demonstrated poor abstract reasoning abilities of LLMs, it is unlikely that current LLMs will be able to compute arbitrary circuits out-of-the-box. However, recent techniques in prompting LLMs have suggested the ability of LLMs to learn from a few prompts (termed in-context learning) Chan et al., 2022, Zhou et al., 2023, Reddy, 2024, Olsson et al., 2022, Zhou et al., 2022, Wei et al., 2022b. In other words, a few question-answer pairs can be shown to the LLM in the context window, followed by the target problem. This approach naturally offers an approach to measure algebraic generalization across circuit divergence metrics: Given a class of question-answer pairs generated from a specific circuit class, how well can a model generalize to a problem generated from a different complexity class (Fig. 7)? Another approach to systematically evaluate the ability of LLMs to compute circuits is through chain-of-thought prompting. Recent theoretical studies have indicated that leveraging chain-of-thought enables LLMs to solve more challenging problems by enhancing their expressivity and using scratchpads Feng et al., 2023, Merrill and Sabharwal, 2024, Li et al., 2024. Practically, inducing chain-of-thought in LLMs provides a way to verify whether LLMs sequentially “reason” through steps that are analogous to ground truth circuit computations (i.e., following the edges from input gates). These experimental approaches provide concrete methods from which to quantify symbolic ability in pretrained LLMs.
@@
Circuits provide a useful formalism to study generalization in AI systems. They are also a leading approach to quantifying the computational complexity of algorithmic problems. However, exactly how formal models of computation are related to AI generalization is broadly unexplored. Here we highlight three primary challenges associated with linking formal models of computation with AI generalization.
@@
Algebraic circuit complexity studies the computational resources (in terms of circuit properties) required to evaluate polynomials. Though there are other metrics of complexity, such as Kolmogorov Complexity in algorithmic information theory Kolmogorov, 1968, Li and Vitányi, 2019, this measure of complexity is incomputable, as it requires searching over an infinite number of programs (though there are efforts in approximating Kolmogorov Complexity with alternative methods; Wyeth and Sturtivant 2023, Johnston et al. 2022, Dingle et al. 2018). Thus, while circuits allow for the explicit computation of a problem, it remains unclear as to whether notions from circuit complexity will elegantly map onto notions of AI generalization. Nevertheless, given the lack of any framework to measure complexity in AI for symbolic ability, we believe that providing a general circuit complexity framework from which to design quantitative benchmarks (via circuit divergence measures) will be an important step towards building a science of symbolic generalization (e.g., see also Dziri et al. 2023, Ram et al. 2023). More broadly, an algebraic circuit approach offers extensive machinery to enable generalization through other algebraic tools, such as minimum spanning/basis sets, their decompositions, and more.
Extending formal models of symbolic computation to natural language and beyond
The types of algebraic problems we presented here are idealized problems that can be straightforwardly mapped to a circuit encoding. However, many real world problems that require algebraic ability are posed in natural language. Thus, applying a circuit complexity-theoretic approach for natural language problems requires parsing natural language into a circuit encoding, adding difficulty to the already challenging problem of generalizing across circuit complexities. It will be important for future work to design frameworks and benchmarks that require tests of algebraic generalization in terms of circuit divergence in natural language settings.
Algebraic representation learning in AI models
@@
The standard approach to training AI models to perform algebraic or compositional tasks is to directly train them with input-output pairs. While algebraic circuits encode a specific algorithm by which to compute that polynomial (i.e., follow the edges from the input gates), it is possible that there are other viable algorithms to compute that polynomial. (For example, it could be possible that an algorithm first factorizes a sum of monomial representation prior to computing that simplified circuit.) More generally, it can be difficult to ascertain the conditions by which a learning system learns a correct and reliable algorithm to solve algebraic problems using only input-output pairs. Naturally, neurosymbolic approaches are often well-suited to learning algebraic and symbolic problems that are algorithmically reliable and sample efficient (e.g., Klinger et al. 2023, Lake et al. 2015, Poesia and Goodman 2023, Trinh et al. 2024). However, designing general purpose neurosymbolic models can be challenging. By contrast, though statistical machine learning models (e.g., transformers) are “general purpose”, the learning process is often obfuscated by learning dynamics that depend on architecture and complex optimization protocols. This makes it difficult to ascertain what computations and algorithms statistical systems learn. However, recent studies in compositional representation learning have suggested that different factors – such as choice of initialization and/or training curriculum – can have a strong influence on whether a model learns compositionally Lippl and Stachenfeld, 2024, Ito et al., 2022, Lake and Baroni, 2023, Zhang et al., 2024, Saglietti et al., 2022. In addition, developing techniques from the field of mechanistic interpretability provide new avenues from which to inspect whether the learned representations in an AI model are faithful to the hypothesized underlying algorithm (e.g., Fig. 6; Olsson et al. 2022, Friedman et al. 2023). Nevertheless, leveraging diverse methods to carefully investigate the algorithms that symbolic and statistical AI models learn will be important for their interpretability, reliability, and overall safety to ensure reliable deployment of AI systems.

### v2

@@ Introduction
Systematic compositional generalization and regression. Systematic compositional generalization refers to the ability to recombine known basis elements into novel combinations of fixed sequence size (Fig 4A). This means that test sets of systematicity are limited to novel combinations of the same length as seen during training 8, 9. The analog in algebraic circuits is to 1) sample a family of circuits $\mathcal{C}$ over a field $\mathbb{F}$ of a fixed size and depth, and 2) to generalize to circuits of the same size and depth, but differentially sampling input gates and/or operators. This can be implemented by choosing different samplers – $P_{1}$ and $P_{2}$ – that differentially sample gates (Fig 4A). For example, the training set of circuits could be constructed with input gates $a_{i}\in_{P_{1}}\mathbb{F}$, and the testing set of circuits could be constructed with input gates $b_{i}\in_{P_{2}}\mathbb{F}$. (Note that when samplers preferentially choose $b_{i}\gg a_{i}$, this is analogous to the common test of “length generalization”; Fig. 2A.)
In real world data, distributional properties of the training distribution can often bias AI (or even human) learners towards learning memorized short cuts, rather than learning algorithmic or syntactic strategies that enable robust generalization7. However, given the infinite vocabulary (i.e., field elements) and the ability to sample circuits with well-defined algorithmic properties, algebraic circuit complexity provides a comprehensive experimental sandbox for designing and sampling from diverse data distributions. Through this sandbox, practitioners can identify specific properties of train/test distributions that can either introduce a distributional bias/confound to be assessed, or ameliorate these biases by counterbalancing the distribution. Interestingly, we note that successful systematic compositional generalization on input gates sampled from different distributions amounts to learning a distributionally robust regression model (e.g., Zhang et al. 55, Ghosh et al. 56). This is because a circuit of specific size and depth (e.g., as shown in Fig. 4A) expresses a specific polynomial. Generalization over a fixed circuit with a distribution shift of its input gates (e.g., field elements but not operator gates) would demonstrate successful out-of-distribution generalization.
Prior notions of systematic compositional generalization have specified “weak” and “strong” forms of systematicity57. “Strong” systematicity refers to the ability to generalize to tokens (e.g., operands) in novel syntactic positions. (For example, if a model is exclusively trained on expressions of the form $a+b$ and $a+c$, where $a$ is always the first token, a model that generalizes to $b+a$ where $a$ is the last token would exhibit strong systematic generalization.) “Weak” systematicity refers to the ability to exclusively generalize to novel expressions in which $a$ is always in the same syntactic position (e.g., always the first token). While this distinction does not alter the semantics for commutative algebras (as discussed in this paper), this distinction is important for many formal languages in which permuting syntactic ordering produces novel semantics, such as in many context-free grammars and non-commutative algebras. (In these cases, circuits can be encoded as ordered DAGs, in which the order of parent/input gates is preserved.)
@@
Figure 4: Analogs of common compositional generalization benchmarks in terms of algebraic circuits. A) A simpler form of generalization within an algebraic circuit is generalizing to circuits of the same structure (size and depth), but with a novel combination of gates. The analog of this in compositionality is commonly-referred to as systematic compositional generalization 9. (Learning over circuits of fixed structure can also be viewed as learning a regression model 55.) On the left, we illustrate an example of a model that is trained on a restricted family of circuits where the input gates are sampled from a field $\mathbb{F}$ with a sampling function $P_{1}$. At test time, the model is required to generalize to circuits of the same circuit class (in terms of size and depth), but where the input gates are sampled using a different sampler $P_{2}$. On the right, both input and operator gates are chosen with separate samplers across train and test circuits, resulting in a more difficult test of systematic compositionality. B) Productive compositional generalization considers partitions of the training and testing sets across circuit sizes and depths. The analog of this in compositionality is commonly-referred to as productive compositional generalization. Given a family of circuits $\mathcal{C}_{\mathbb{F},X,D,S}$, where $D$ and $S$ denote a set of depths $d_{i}$ and size $s_{i}$, the primary experimental manipulation is to construct a training set $\mathcal{C}_{\mathbb{F},X,D_{1},S_{1}}$ and testing set $\mathcal{C}_{\mathbb{F},X,D_{2},S_{2}}$ such that there exists no overlap of a specific circuit class $C_{\mathbb{F},X,d_{i},s_{i}}$ between the training and testing sets.
@@
Productive compositional generalization. Productive compositional generalization refers to the ability to generalize to sequences of longer length. In the context of algebraic circuits, while systematicity focuses on keeping circuit size and depth fixed while manipulating gates, productive generalization focuses on manipulating circuit size and depth (Fig 4B). Thus, evaluating systematic and productive generalization can be studied together; measures of generalization can be quantified in a continuous manner by varying circuit parameters, such as gate samplers (systematicity), size (productivity), and depth (productivity). Understanding how each of these properties interact across training and testing sets will provide a comprehensive quantification of generalization over distinct algorithmic properties.
@@
Since algebraic circuits are circuit representations of algebraic expressions, one can ask more generic questions about algebraic polynomials. For example: Given a class of polynomials as a training dataset, what other class of polynomials will this model be able to compute? Such a question goes beyond asking whether a model can systematically or productively generalize. Instead, it addresses a basic question that can leverage other rich mathematical subfields (e.g., geometry, topology) to quantify algorithmic generalization. Algebraic circuits provide a flexible framework to formalize problem complexity beyond existing paradigms in compositional generalization.
@@
Prior work studying arithmetic abilities in transformer models have typically focused on computing simple expressions with field elements ($a\in\mathbb{F}$; e.g., $5+7=?$), rather than abstract variables ($x\in X$; e.g., $72x_{1}+x_{2}+$ 7). Including variables in an algebraic expression increases the polynomial degree ($d\geq 1$), thereby increasing its complexity and the need for abstractions. One approach to evaluating algebraic circuits with variables is through polynomial identity testing, an active area of research in computational complexity theory and computational algebra. Polynomial identity testing evaluates whether two polynomials are equivalent (i.e., $P_{1}(x_{1},...,x_{n})\equiv P_{2}(x_{1},...,x_{n})$). Posed another way, one can ask whether $P_{1}(x_{1},...,x_{n})-P_{2}(x_{1},...,x_{n})\equiv 0$.
@@
This problem can be naturally formulated as a binary classification task for AI models. Importantly, studying the extent of algebraic generalization by assessing train/test circuit divergences of polynomial identity problems can be rigorously quantified, since each identity problem can be encoded as a circuit (Fig. 5A). The complexity of a polynomial identity testing problem can be scaled up by increasing the size, depth, and/or degree of the test problems. Moreover, machine learning practitioners can leverage existing symbolic programs and solvers to systematically generate large datasets (e.g., SymPy; Meurer et al. 58). Polynomial identity testing thus provides a unique opportunity to study AI generalization in terms of circuit complexity theory.
@@
Figure 5: Algebraic problems as machine learning challenges. Prior work that leverages arithmetic problems for machine learning studies are typically limited to evaluating expressions with field elements. We introduce problems that can be evaluated with abstract variables. A) Polynomial identity testing as a machine learning classification task. Polynomial identity testing, an important and active area of research in computational algebra, evaluates whether two polynomials are equivalent. We illustrate two different polynomial expressions, $P_{1}(x)$ and $P_{2}(x)$ with distinct circuit representations, yet are mathematically equivalent. To reformulate this as a classification task for machine learning studies, one can ask if $P_{2}(x)-P_{1}(x)\equiv 0$ (right). B) Polynomial expansion and/or factorization as a transduction (sequence-to-sequence) task. Common sequence-to-sequence tasks in linguistics ask whether a model can expand a string using a set of rules or a grammar, such as SCAN or PCFG 8, 9. In algebraic circuits, this is analogous to expanding a polynomial in a factorized representation (left). Given the 1-1 correspondence of polynomials, an additional approach is to take a polynomial in its expanded form (sum of monomials), and generate the factorized representation (right). This provides the ability to evaluate whether an AI system can expand an encoding (expansion), or compress an encoding (factorization).
@@
A wide application of generative models is in sequence-to-sequence transduction tasks. One particularly impactful use case is the development of AI models for code. AI models for code take in code as inputs (e.g., COBOL), and generate a translation of that code (e.g., Java). Despite its potential importance for modernizing many existing codebases, there is significant skepticism as to whether generative models trained with next-token prediction can reliably generate accurate code translations. This is due to the fact that programming languages are not dictated by autoregressive processes, and instead governed by algorithmic rules.
Learning on algebraic circuits provides a straightforward framework to evaluate the ability of models to learn sequence-to-sequence tasks that are governed by algorithmic rules. For example, the problem of expanding or factorizing a polynomial is a problem that is governed by the axiomatic rules of algebra (Fig. 5B). Like code translation, this task requires transforming one sequence into another while maintaining mathematical equivalence (or for code, semantic equivalence despite syntactic differences). Importantly, there are explicit tools to describe the complexity with which the translation occurs in algebraic circuits. For example, a polynomial expansion transforms a factorized representation (i.e., shallow circuit) to a sum of monomials representation (i.e., deeper circuit). In contrast, a polynomial factorization implements the inverse operation, which requires compressing a large circuit to a smaller circuit. While a general interesting question is to ask whether models trained on one type of transformation can learn the other, this formalization has natural implications for understanding how to design AI models for code. In particular, some programming languages may have lower-level syntax (e.g., COBOL) relative to other languages, such as Python or Java. In either case, understanding how lower-level and higher-level encodings of an expression (be they algebraic or programmatic) requires learning useful abstractions and the algorithms to translate them. Thus, studying the conditions by which AI models can robustly parse and translate algebraic circuits can shed light on the best strategies to train AI models to translate programming languages.
@@
A major issue in assessing algorithmic generalization is the difficulty of interpreting what goes awry when they fail to generalize. This is partly due to the lack of interpretability of many benchmarks, which are often presented in natural language, and where verifiable algorithms (e.g., circuit diagrams or parse trees) do not exist 14, 59. A more recent set of approaches in characterizing the interpretability of neural network representations rely on the design of carefully constructed tasks that have interpretable tasks, such as ground truth parse trees or circuit diagrams 60, 24, 36. Algebraic circuits similarly provide verifiable tasks and sub-tasks. While the input to an AI model might be a string representation of a polynomial, that polynomial’s circuit encoding provides its ground truth encoding. Such an encoding makes it easy to directly compare to the internal representations of a model, such as the attention weights within a transformer (Fig. 6). This makes it possible to track if the internal representations of a transformer computes the algebraic expression similarly to the algorithm encoded by its circuit – the normative algorithm to computing polynomials. More broadly, the distance metric that computes the distance between attention weights and ground truth circuit edges (as shown in Fig. 6) can be used as a regularization term to encourage transformers to learn normative circuit computations, providing algorithmic and step-by-step information to the model. Overall, evaluating AI models on algebraic problems allows practitioners to adjudicate between theoretical models of circuit computation with modern AI computation.
Figure 6: Using an algebraic circuit’s adjacency matrix as a ground truth comparison to interpret transformer attention representations. The transformer’s attention mechanism provides a useful way to peer into the representations of its input tokens (Left). When the input is an algebraic expression (presented as a string of tokens), the attention matrix can be investigated to uncover the relationships between tokens (i.e., operators and operands). On the other hand, an algebraic expression can always be mapped to a circuit encoding, which can be represented as the adjacency matrix of that circuit (Right). This allows for the direct comparison between attention weights and the ground truth circuit representation (Middle). This distance between the transformer’s attention weights and the ground truth circuit adjacency matrix can also be used as a regularizer to encourage learning circuit algorithms via attention weights.
@@
We have introduced algebraic circuit tasks within the context of training models from scratch. While it is difficult to evaluate pretrained LLMs with precise levels of certainty due to the obscure nature of the pretraining data and optimization protocols, algebraic tasks can still be repurposed for LLMs. Prior work has demonstrated that LLMs typically fail on mathematical problems which require abstract reasoning on variables 61, 25, 7. Thus, if an LLM calls specific tools (e.g., calculators) which make computing circuits with only field elements too simple (e.g., those illustrated in Fig. 1 and 2, we can replace these with circuits containing variables $x\in X$ (i.e., those of degree $\geq 1$)). For example, such tasks can include the aforementioned polynomial identity testing or polynomial expansion/factorization tasks, both of which can be made arbitrarily difficult and use arbitrary sets of variables/tokens.
Given that prior studies have demonstrated poor abstract reasoning abilities of LLMs, it is unlikely that current LLMs will be able to compute arbitrary circuits out-of-the-box. However, recent techniques in prompting LLMs have suggested the ability of LLMs to learn from a few prompts (termed in-context learning) 62, 37, 63, 64, 65, 66, 67. In other words, a few question-answer pairs can be shown to the LLM in the context window, followed by the target problem. This approach naturally offers an approach to measure algorithmic generalization across circuit divergence metrics: Given a class of question-answer pairs generated from a specific circuit class, how well can a model generalize to a problem generated from a different complexity class (Fig. 7)? Another approach to systematically evaluate the ability of LLMs to compute circuits is through chain-of-thought prompting. Recent theoretical studies have indicated that leveraging chain-of-thought enables LLMs to solve more challenging problems by enhancing their expressivity and using scratchpads 68, 48, 69, 70. Practically, inducing chain-of-thought in LLMs provides a way to verify whether LLMs sequentially “reason” through steps that are analogous to ground truth circuit computations (i.e., following the edges from input gates). Moreover, since algebraic circuits encode an explicit algorithm to compute an expression, each step through the circuit produces a verifiable intermediate computation. This enables the evaluation of LLMs not only on verified outputs, but also verified intermediate reasoning chains, providing an opportunity to evaluate reinforcement learning approaches that have been increasingly used to develop reasoning-based models 71, 4 Together, these experimental approaches provide concrete methods from which to quantify algorithmic ability in LLMs.
@@
Circuits provide a useful formalism to study algorithmic generalization in AI systems. They are also a leading approach to quantifying the computational complexity of algorithmic problems in theoretical computer science. Prior AI reasoning literature has focused on benchmarking AI to human-level reasoning72, 73, 74, 75, or on how humans learn algorithmic tasks76, 77, 49, both of which are important areas of investigation from the perspectives of cognitive science and AI. However, the present Perspective provides a complementary viewpoint from theoretical computer science aimed to quantify reasoning based on generalizing over algorithmic complexity. Here we highlight the primary challenges associated with linking formal circuit models of computation with AI generalization.
@@
Algebraic circuit complexity studies the algorithmic resources required to evaluate polynomials. Though there are other metrics of complexity, such as Kolmogorov Complexity in algorithmic information theory 78, 79, this measure of complexity is incomputable, as it requires searching over an infinite number of programs (though there are efforts in approximating Kolmogorov Complexity with alternative methods; Wyeth and Sturtivant 45, Johnston et al. 80, Dingle et al. 81). Thus, while circuits allow for the explicit computation of a problem, it remains unclear as to whether notions from circuit complexity will naturally map onto notions of AI generalization.
Other studies have investigated the theoretical requirements for transformer models to evaluate formal languages and algorithms69, 48, 82, 83, 84. While these studies provide useful insight into what architectural components are likely important to implement algorithmic problems of a particular complexity (e.g., deeper networks for circuit depth and wider context windows for circuit size), these studies do not address the learnability of circuit algorithms. Furthermore, it will be interesting to understand how frameworks grounded in algorithmic circuit complexity (as presented here) relate to other measures of compositionality, which are often measured by properties of the computation graph of a task24, divergence of input-output mapping of a task46, or divergence of a distribution of training and testing tasks (e.g., MCD)17.
Nevertheless, given the lack of any framework to measure algorithmic complexity in AI, we believe that introducing a circuit complexity framework from which to design quantitative benchmarks will be an important step towards building a science of algorithmic generalization. Furthermore, an algebraic circuit approach offers extensive machinery to enable generalization through other algebraic tools, such as minimum spanning/basis sets, their decompositions, and more.
Faithful algorithmic representation learning
Though an algebraic circuit encodes an explicit algorithm to compute a polynomial at a particular level of abstraction (i.e., follow the edges from the input gates), it is possible that there are alternative viable algorithms to compute that same polynomial. For example, would an AI model simply follow the edges from the input gates to the output gates? Or might it factorize the expression (leading to a shallower circuit) prior to evaluating that expression? Relatedly, the algorithmic steps required to implement long addition (particularly when computing the sum of two very large numbers) are not fully captured in the algebraic circuits we present in this Perspective. (For example, long addition as specified in Dziri et al. 24 requires additional operators such as “carry”, “concatenate”, “modulo”, etc.) Would AI models implement long addition with a different set of computational gates and operators to accommodate arithmetic with large numbers? Nevertheless, despite these potential ambiguities, using a computational circuit framework provides testable and verifiable hypotheses that allow us to empirically evaluate what algorithm a model implements. Furthermore, use of a circuit framework enables the design of quantitatively meaningful algorithmic benchmarks such as those designed to test generalization over algorithmic time complexity (circuit depth), or space complexity (circuit size), among others. More broadly, a computational circuit framework can naturally extend beyond algebraic circuits to formal languages, such as those within the Chomsky hierarchy47 (e.g., context-free grammars), by characterizing the circuit parameters of the language’s parse tree. Thus, this framework provides a unified theoretical foundation for quantifying algorithmic complexity across diverse empirical evaluations, including arithmetic tasks, formal languages, and other compositional paradigms, and is an important step toward understanding the faithfulness by which an AI system computes a class of problems.
@@
Alternatively, many prior approaches to faithfully learning algorithmic representations often involve neurosymbolic methods. These methods provide promising avenues to learn discrete algorithmic solutions to problems that are reliable and sample efficient53, 85, 86, 87, 88. However, designing general purpose (rather than domain-specific) neurosymbolic models can be challenging, since they are often not fully differentiable or require strong inductive biases. By contrast, though statistical machine learning models (e.g., transformers) are “general purpose”, the learning process is often obfuscated by learning dynamics that depend on architecture and complex optimization protocols. This makes it difficult to ascertain what algorithms statistical systems learn. However, recent studies in compositional representation learning have suggested that different factors – such as choice of initialization and/or training curriculum – can have a strong influence on whether a model learns compositionally 89, 90, 49, 91, 92. In addition, developing techniques from the field of mechanistic interpretability provide new avenues from which to inspect whether the learned representations in an AI model are faithful to the hypothesized underlying algorithm (e.g., Fig. 6; Olsson et al. 64, Friedman et al. 93). Nevertheless, leveraging diverse methods to carefully investigate the algorithms that symbolic and statistical AI models learn will be important for their interpretability, reliability, and overall safety to ensure reliable deployment of AI systems.